sound of frequency 600 Hz fall normally on a perfectly reflecting wall. The shortest distance from the wall at which the air particles have the maximum amplitude of vibration is

Grade:11

1 Answers

Arun
6 years ago
Dear Garv
Shortest distance should be lambda/4
 
Now
Lambda = v/f = 330/600 = 0.55
 
 
Hence distance = 0.55/4 = 0.1375 m
 
Regards
Arun (askIITians forum expert)